OpenStudy (anonymous):

i need help. how to solve: -3c - 12 = -5 + c

OpenStudy (karatechopper):

c=-7/4

OpenStudy (anonymous):

can you also show me how u did it? :)

OpenStudy (karatechopper):

sure hold on

OpenStudy (anonymous):

okayy, thank you!

OpenStudy (karatechopper):

-3c-12=-5+c -3c+3c-12=-5+c+3c -12=-5+4c -12+5=-5+5+4c -7=4c -7/4=4c/4 -7/4=c

OpenStudy (karatechopper):

hope u understand tht???^

OpenStudy (anonymous):

thank you!! i do now :)

OpenStudy (karatechopper):

good!!